# RoboMM: All-in-One Multimodal Large Model for Robotic Manipulation

### v1

@@ 5 Experiments
In the previous sections, we elaborate on the RoboMM framework and the characteristics of the RoboData dataset. Next, we address the following questions using task success rates: its performance across multiple datasets, the importance of each module within RoboMM.
@@ 5.1 Results and Analysis
Protocol. Due to the substantial time and computational resources required for model training, this study chooses to train the RoboMM models without Image Loss on five datasets: CALVIN [41], Meta-World [65], LIBERO [34], RoboCasa [47], and Robomimic [39], as shown in the Figure 8. A total of 10 epochs are trained, utilizing 2.1 million samples, employing 32 A100 GPUs with 80GB each, and taking 130 hours.
During the evaluation phase, we adhere to the official configurations provided for each dataset. Specifically, for the CALVIN dataset [41], a sequence length of 1000 is used, while for the other datasets, each task was evaluate over 20 episodes with varying initial conditions. The evaluation metric adopted is success rate(SR), to comprehensively assess the model’s performance.
Experiment. The results summarized in Table 1 reveal that RoboMM exhibits exceptional performance across the evaluated datasets, achieving state-of-the-art results in particular on the LIBERO [34] and RoboCasa [47] datasets. Notably, in the CALVIN dataset [41], RoboMM attain a success rate (SR) of 91.0%, which is competitive with the performance of MDT [55] and HULC++ [43], significantly exceeding the performance of models in the second tier. In the Meta-World dataset [65], RoboMM achieve a success rate of 78.6%, positioning it within the top tier of models. Furthermore, RoboMM consistently outperformed other models in LIBERO [34], RoboCasa [47], and Robomimic [39], demonstrating its robustness and adaptability across a variety of tasks and environments. These findings not only underscore RoboMM’s versatility but also highlight its effectiveness in handling tasks with varying levels of complexity, indicating its potential for real-world applications.
In contrast, RoboMM- reflects the results obtained from training on unaligned RoboData. The findings indicate that the misalignment between input and output spaces significantly hampers model performance when training across multiple datasets. It is important to note that in the Meta-World dataset [65], the impact of this misalignment is relatively minimal. This is because the actions involved are limited to three positional changes without rotation ($a_{pose}^{h}=\{\Delta\text{pos}_{x}^{h},\Delta\text{pos}_{y}^{h},\Delta\text{pos}_{z}^{h},0.0,0.0,0.0\}$), and the original spatial coordinate system aligns with that of the RoboData coordinate system. As a result, there is no need for additional alignment processes, allowing the model to perform adequately despite the lack of training on aligned data.
The successful alignment of input and output spaces within RoboData, combined with the compatible architecture of RoboMM, enables effective joint training and evaluation across multiple datasets. It is important to note that, unlike RoboMM, a Generalist Policy, all other compared models are specialized and limited to training or fine-tuning on a single dataset. While many studies in the field may involve training on multiple datasets, they often necessitate fine-tuning on the final evaluation dataset. This domain-specific fine-tuning can enhance performance but also increases training costs and may compromise the generalization capabilities of the models.
5.2 Ablation Study
Protocol. This experiment aims to systematically evaluate the contributions of each module in the RoboMM framework to overall performance, using the CALVIN [41] dataset for analysis. We design six different experimental setups to progressively analyze the impact of each module. First, in the baseline setup, the model receives $H$ frames as input and outputs an action only at the last frame to assess its basic prediction capability. Next, with the addition of the FFA module, the model is able to output actions after each frame, allowing us to examine its immediate prediction capability and performance in continuous prediction scenarios. As the experiment progresses, the introduction of the Image module enables the model to generate the next frame, thereby enhancing its contextual understanding and fine-grained perception. Subsequently, we added the UVFormer module to assess its specific contribution to the model’s performance. Finally, by incorporating the occupancy (OCC) output, we further enhance the model’s understanding of 3D spatial information, thereby improving its predictive capabilities.
@@ 5.2 Ablation Study
Table 2: Performance on different modules of RoboMM. FFA and OCC refer to frame by frame action module and occupancy supervision.
Experiment. The results of the experiments are summarized in Table 2, which demonstrates a clear trend of performance improvement as different modules are added. The baseline configuration achieves a success rate of 81.0% for completing the first task, while the addition of the Frame by Frame Action (FFA) module increases this success rate to 85.0%. The introduction of image output capability further enhances success rates across all tasks. Notably, the UVFormer setup shows a significant improvement in the first task, achieving a success rate of 94.2%, indicating the effectiveness of the UVFormer module in enhancing the model’s predictive capabilities. The OCC configuration builds on this by incorporating occupancy loss, resulting in even higher success rates, particularly in later tasks, suggesting that the model benefits from enriched contextual and spatial information.
Finally, row 6 shows the highest performance across all tasks, with a significant increase in the average length of completed tasks. Combining insights from Table 2 and Figure 7, it is evident that although the generated images and occupancy do not exhibit ideal quality, they significantly contribute to enhancing action performance. This indicates that even with suboptimal generation quality, the integration of visual and spatial information still has a positive impact on the execution of robotic tasks.
Overall, this ablation study highlights the contributions of each module within RoboMM, illustrating how enhancements can lead to significant performance gains in task completion.
5.3 Comparison with OpenVLA
To evaluate the superiority of our model architecture, we compare RoboMM with the currently best-performing OpenVLA. To ensure a fair comparison, we set the window size to 1, and train RoboMM from scratch, while OpenVLA is fine-tuned on the officially released weights.
As shown in Table 3, RoboMM outperforms OpenVLA (LoRA) in multiple metrics, particularly in Task 3-5 and average sequence length. This indicates that RoboMM can capture longer dependencies when handling tasks, thereby improving model accuracy. These results not only demonstrate the superior performance of the RoboMM architecture but also provide valuable references for future research.
@@ 5.3 Comparison with OpenVLA
Figure 7: The illustration demonstrating the auxiliary modality results of RoboMM on the Calvin three tasks. The inputs consist of static images, wrist-view images, and text, while the outputs include the current OCC and the predicted next frame’s static and wrist-view images.
Task Completed in a Sequence Avg Len 1 2 3 4 5 OpenVLA (LoRA) 78% 55% 29% 17% 8% 1.86 RoboMM (ours) 81% 54% 37% 25% 16% 2.15
Table 3: Performance comparison with OpenVLA on CALVIN.
6 Conclusion
This paper presents RoboMM and RoboData, two innovative solutions designed to address key challenges in robotic learning. We have significantly improved the performance of robotic manipulation tasks by integrating multiple datasets and utilizing advanced model architectures.
RoboMM enhances 3D spatial perception through the introduction of camera parameter calibration and occupancy supervision, while also improving fine-grained perception capabilities via multimodal output. In parallel, RoboData integrates multiple industry datasets, supplements missing modalities such as depth and camera parameters, and ensures proper alignment of input and output spaces for better model performance.
@@ 6 Conclusion
Although our training objectives in supervised learning have reached only a moderate scale, RoboMM and RoboData open new avenues for cross-embodiment joint training and evaluation. We hope this work provides the industry with a comprehensive and fair evaluation system, inspires future exploration in foundational robotic model research, and enhances the generality and performance of robotic learning overall.
@@
7 RoboMM detailed information
@@ 9.2 Meta-World Dataset
Figure 8: Evaluation Datasets. We evaluate RoboMM across five simulation benchmarks and present policy rollout visualizations of the experiments. From left to right, the benchmarks include CALVIN, Meta-World, LIBERO, RoboCasa, and Robomimic. Experiment details can be found in Section 5.1.
@@ 10 Experiment Details
The success rates of the expert models in the Table 1 are organized from the following sources: The evaluation methods on the CALVIN [41] dataset are sourced from the official CALVIN leaderboard (url: http://calvin.cs.uni-freiburg.de/). In the Meta-World [65] dataset, the results of PAD [23], GR-1 [64], SuSIE [5], RT-2* [7], and RT-1 [6] come from PAD [23], while the results of PRISE [69] are derived from related papers. In the Libero [34] dataset, the results of QueST [45], VQ-BeT [31], PRISE [69], DiffusionPolicy [15], ACT [67], and ResNet-T [34] all come from QueST [45], while the results of MDT [55] and Distill-D [24] are sourced from MDT [55]; the results of MaIL [27], ATM [63], and MUTEX [56] come from their respective research papers. The results of RoboCasa [48] in the RoboCasa [48] dataset are sourced from related papers. In the Robomimic [39] dataset, the results of IBC [19] come from personal replication.
The success rates of each task for Robomimic on various datasets are shown in the Table 5,  6,  7,  8,  9.
Task Success Rate
Rotate Blue Block Right 87.9%
Move Slider Right 86.8%
Turn Off LED 87.7%
Push Into Drawer 45.2%
Lift Blue Block Drawer 93.3%
Lift Pink Block Slider 56.2%
Place In Slider 43.4%
Open Drawer 90.1%
Rotate Red Block Right 81.2%
Lift Pink Block Table 82.9%
Push Blue Block Left 88.3%
Close Drawer 86.9%
Turn On LED 92.4%
Stack Block 39.2%
Push Pink Block Right 75.4%
Rotate Blue Block Left 89.7%
Lift Blue Block Table 86.7%
Place In Drawer 86.2%
Turn On Lightbulb 81.9%
Move Slider Left 84.8%
Rotate Red Block Left 90.0%
Lift Red Block Slider 62.5%
Push Pink Block Left 86.4%
Push Red Block Left 70.0%
Lift Blue Block Slider 64.8%
Push Red Block Right 69.7%
Lift Red Block Table 82.3%
Turn Off Lightbulb 90.7%
Lift Pink Block Drawer 70.0%
Rotate Pink Block Right 56.7%
Rotate Pink Block Left 83.3%
Push Blue Block Right 41.0%
Unstack Block 78.9%
Lift Red Block Drawer 85.7%
Table 5: RoboMM Success Rates on Various Tasks in CALVIN [41].
Task Success Rate (%)
Assembly 65%
Basketball 100%
Bin Picking 75%
Box Close 75%
Button Press Topdown 100%
Button Press Topdown Wall 100%
Button Press 100%
Button Press Wall 90%
Coffee Button 70%
Coffee Pull 40%
Coffee Push 70%
Dial Turn 90%
Disassemble 50%
Door Close 100%
Door Lock 100%
Door Open 100%
Door Unlock 100%
Hand Insert 70%
Drawer Close 100%
Drawer Open 100%
Faucet Open 0%
Faucet Close 100%
Hammer 35%
Handle Press Side 100%
Handle Press 100%
Handle Pull Side 20%
Handle Pull 50%
Lever Pull 100%
Peg Insert Side 65%
Pick Place Wall 100%
Pick Out of Hole 50%
Reach 50%
Push Back 80%
Push 95%
Pick Place 85%
Plate Slide 100%
Plate Slide Side 100%
Plate Slide Back 100%
Plate Slide Back Side 100%
Peg Unplug Side 25%
Soccer 45%
Stick Push 100%
Stick Pull 85%
Push Wall 95%
Reach Wall 90%
Shelf Place 20%
Sweep Into 95%
Sweep 65%
Window Open 85%
Window Close 100%
Table 6: RoboMM Success Rates on Various Tasks in Meta-World [65]
Task Success Rate
Lift 45%
Can 0
Square 0
Table 7: RoboMM Success Rates on Various Tasks in RoboMimic [39].
Task Index Success Rate (%) Task Index Success Rate (%)
0 100% 1 40%
2 100% 3 95%
4 95% 5 100%
6 85% 7 100%
8 100% 9 100%
10 100% 11 100%
12 75% 13 95%
14 100% 15 100%
16 90% 17 85%
18 100% 19 85%
20 100% 21 100%
22 100% 23 55%
24 100% 25 95%
26 90% 27 95%
28 100% 29 100%
30 65% 31 90%
32 30% 33 75%
34 100% 35 100%
36 90% 37 100%
38 90% 39 100%
40 80% 41 100%
42 100% 43 85%
44 100% 45 100%
46 100% 47 100%
48 95% 49 100%
50 100% 51 20%
52 100% 53 85%
54 100% 55 100%
56 100% 57 100%
58 100% 59 100%
60 100% 61 95%
62 95% 63 100%
64 70% 65 100%
66 100% 67 85%
68 100% 69 100%
70 100% 71 100%
72 90% 73 75%
74 100% 75 55%
76 90% 77 95%
78 90% 79 100%
80 75% 81 25%
82 100% 83 80%
84 90% 85 75%
86 100% 87 100%
88 100% 89 95%
Table 8: RoboMM Success Rates on Various Tasks in Libero [34].
Task Name Success Rate (%)
CoffeePressButton 90%
CoffeeServeMug 55%
@@
CloseDrawer 95%
OpenDrawer 50%
TurnOffMicrowave 35%
TurnOnMicrowave 70%
PnPCabToCounter 30%
PnPCounterToCab 30%
PnPCounterToMicrowave 15%
PnPCounterToSink 30%
PnPCounterToStove 30%
PnPMicrowaveToCounter 30%
PnPSinkToCounter 5%
PnPStoveToCounter 20%
@@
Table 9: RoboMM Success Rates on Various Tasks in RoboCasa [48].

### v2

@@ 4 Experiments
In the previous sections, we elaborate on the RoboTron-Mani framework and the characteristics of the RoboData dataset. Next, we address the following questions using task success rates: its performance across multiple datasets, the importance of each module within RoboTron-Mani and RoboData.
@@ 4.1 Results and Analysis
Table 2: Performance on different modules of RoboTron-Mani. FFA and OCC refer to frame by frame action module and occupancy supervision. Row 7* signifies the replacement of the MLP action head with the DiT action head.
Protocol. Due to the substantial time and computational resources required for model training, this study conducts joint training of the RoboTron-Mani model with DiT action head across five datasets. These datasets encompass simulated data (CALVIN [43], Meta-World [68], LIBERO [36], RoboCasa [49]) and RT-1 [8], as shown in the Figure 9. The RoboTron-Mani model, with 4 billion parameters in bf16 precision, is trained for 10 epochs using 2.1 million samples on 32 80G-A100 GPUs, taking about 50 hours. During the evaluation phase, we adhere to the official configurations provided for each dataset. The evaluation metric adopted is success rate(SR), to comprehensively assess the model’s performance.
In this study, we are conducting systematic experiments to evaluate the performance of the RoboTron-Mani model in multi-data fusion training and its comparison with expert models. As shown in Table 1, RoboTron-Mani demonstrates impressive performance across several simulation datasets. In the LIBERO [36] dataset, RoboTron-Mani achieves a success rate of 91.7%, significantly surpassing the current best model, QueST [47], which has an 89.8% success rate. In the RoboCasa [49] dataset, the model achieves a success rate of 47.4%, showcasing its robust adaptability in diverse scenarios. For the CALVIN [43] and Meta-World [68] datasets, the success rates are 93.8% and 80.1%, respectively, comparable to other leading expert models such as the MDT [58] and the PRISE [72] model. This demonstrates its balanced capabilities in temporal reasoning and multi-task learning.
Notably, RoboTron-Mani outperforms the general model RoboFlamingo [35] across all four datasets. For instance, there is a 19.6% improvement in success rate in the LIBERO [36] dataset and a 14.8% increase in the Meta-World [68] dataset. This significant performance gap not only highlights the architectural advantages and innovation of RoboTron-Mani but also reflects the potential and superiority of its spatial awareness and multimodal integration strategies in executing complex tasks.
Specifically, as shown in Table 1, RoboTron-Mani achieves a 60% average task success rate on the Google RT-1 [8] benchmark dataset, significantly outperforming models with equivalent parameter sizes. The research focuses on extracting three complex operational tasks (Pick Coke Can, Move Near, Open/Close Drawer). Through structured reassembly and distribution alignment techniques, we construct the multimodal joint training framework, which combines with four RoboData subsets. The system validation is completed under the SimpleEnv evaluation framework. In terms of experimental design, to address the sensor parameter missing issues in the RT-1 [8] dataset, we supplement them based on the SimpleEnv test environment. Given the inaccessibility of depth perception data, we omit occlusion-aware loss during training. Moreover, to handle the lack of wrist camera sensor data, we introduce a dynamic mask filling mechanism during the preprocessing stage.
It is important to note that, unlike the general representation learning paradigm adopted by RoboTron-Mani and RoboFlamingo, all compared models are expert models utilizing task-specific optimization strategies, which restrict parameter updates to a single data domain. Although existing studies include multi-dataset pretraining schemes, they typically rely on fine-tuning adaptations to the target dataset. This domain-dependent optimization approach can achieve local performance gains but also increases training complexity and significantly reduces the model’s cross-domain generalization ability.
4.2 Ablation Study
@@
Figure 7: The illustration demonstrating the auxiliary modality results of RoboTron-Mani on the Calvin three tasks. The inputs consist of static images, wrist-view images, and text, while the outputs include the current OCC and the predicted next frame’s static and wrist-view images.
4.2.1 RoboTron-Mani Module
Protocol. This experiment aims to systematically evaluate the contributions of each module in the RoboTron-Mani framework to overall performance, using the CALVIN [43] dataset for analysis. We design seven different experimental setups to progressively analyze the impact of each module. First, in the baseline setup, the model receives $H$ frames as input and outputs an action only at the last frame to assess its basic prediction capability. Next, with the addition of the Frame by Frame Action (FFA) module, the model is able to output actions after each frame, allowing us to examine its immediate prediction capability and performance in continuous prediction scenarios. As the experiment progresses, the introduction of the Image module enables the model to generate the next frame, thereby enhancing its contextual understanding and fine-grained perception. Subsequently, we added the UVFormer module to assess its specific contribution to the model’s performance. Finally, by incorporating the occupancy (OCC) output, we further enhance the model’s understanding of 3D spatial information, thereby improving its predictive capabilities.
Experiment. The results of the experiments are summarized in Table 2, which demonstrates a clear trend of performance improvement as different modules are added. The baseline configuration achieves a success rate of 81.0% for completing the first task, while the addition of the FFA module increases this success rate to 85.0%. The introduction of image output capability further enhances success rates across all tasks. Notably, the UVFormer setup shows a significant improvement in the first task, achieving a success rate of 94.2%, indicating the effectiveness of the UVFormer module in enhancing the model’s predictive capabilities. The OCC configuration builds on this by incorporating occupancy loss, resulting in even higher success rates, particularly in later tasks, suggesting that the model benefits from enriched contextual and spatial information.
Finally, row 7 shows the highest performance across all tasks, with a significant increase in the average length of completed tasks. Combining insights from Table 2 and Figure 7, it is evident that although the generated images and occupancy do not exhibit ideal quality, they significantly contribute to enhancing action performance. This indicates that even with suboptimal generation quality, the integration of visual and spatial information still has a positive impact on the execution of robotic tasks.
Overall, this ablation study highlights the contributions of each module within RoboTron-Mani, illustrating how enhancements can lead to significant performance gains.
4.2.2 RoboData Alignment
The training of RoboFlamingo and RoboTron-Mani is conducted using both non-aligned and aligned data across the four datasets: LIBERO, RoboCasa, CALVIN, and Meta-World. As indicated in the table, due to RoboFlamingo’s two-dimensional feature input, it cannot achieve input space alignment, resulting in performance that does not meet expectations even when trained on aligned RoboData. The results of RoboTron-Mani- reflect the effects of training on non-aligned RoboData. The research indicates that misalignment between input and output spaces significantly hinders model performance when training across multiple datasets. Notably, in the Meta-World dataset [68], the impact of this misalignment is relatively minimal. This is because the operations are limited to changes in three positions without rotational variations ($a_{pose}^{h}=\{\Delta\text{pos}_{x}^{h},\Delta\text{pos}_{y}^{h},\Delta\text{pos}_{z}^{h},0.0,0.0,0.0\}$), and the original spatial coordinate system is consistent with that of the RoboData. Therefore, no additional alignment process is required, and model performs relatively well even when not trained on aligned data. The successful alignment of input and output spaces within RoboData, combined with the compatible architecture of RoboTron-Mani, enables effective joint training and evaluation across multiple datasets.
Model LIBERO RoboCasa CALVIN Meta-World
RoboFlamingo∗- 69.1% 15.8% 41.7% 63.2%
RoboFlamingo 72.1% 15.6% 43.8% 65.3%
RoboTron-Mani- 64.2% 27.0% 74.7% 79.3%
RoboTron-Mani 90.7% 30.6% 91.0% 78.6%
Table 3: The results of RoboFlamingo and RoboTron-Mani before and after data alignment.
5 Conclusion and Outlook
Enhancing large model performance relies on effectively utilizing vast, consistently formatted data. In the EAI field, robotic operation data is limited. A feasible approach is for institutions to collect data using their robotic devices and consolidate it into a large, heterogeneous dataset to drive industry development. Efficiently using this data is challenging. In RT-X Table 1, direct data fusion results in RT-1-X underperforming compared to RT-1, highlighting the difficulty in utilizing heterogeneous data effectively.
This paper introduces the method to effectively utilize heterogeneous data from various institutions through the model architecture featuring 3D perception input and multimodal output, combined with data reorganization. RoboTron-Mani and RoboData synergize effectively. For instance, capturing the same scene with cameras of different parameters can result in varying 2D visual features, impacting action outputs. In contrast, the 3D scene remains consistent, and UVFormer’s 3D features exhibit minimal variation. Without RoboTron-Mani’s 3D spatial input, achieving input space alignment is impossible. We initially validate this approach using simulations, leveraging the accessibility of heterogeneous data in simulators.
@@ 5 Conclusion and Outlook
Although our training objectives in supervised learning have reached only a moderate scale, RoboTron-Mani and RoboData open new avenues for cross-embodiment joint training and evaluation. We hope this work provides the industry with a comprehensive and fair evaluation system, inspires future exploration in foundational robotic model research, and enhances the generality and performance of robotic learning overall.
@@
6 Related Work
7 Real world experiment
To evaluate RoboTron-Mani performance in real-world scenarios, we constructed a physical evaluation system as illustrated in Figure 8. The robotic platform comprises a Dalu mobile base and an UR3 robotic arm. During experiments, the mobile base remains stationary, and only the robotic arm retaining degrees of freedom. The system is equipped with essential perception and actuation components, including a Robotiq two-finger gripper, an Intel D435 depth camera mounted on the wrist of the arm (denoted as $cam_{wrist}$), and an Orbbec Gemini Pro depth camera fixed to a stand on the ground to the left of the arm (denoted as $cam_{static}$).
Since RoboTron-Mani requires camera extrinsic parameters, we perform hand-eye calibration prior to experiments. Using the image data and intrinsic/extrinsic parameters from both $cam_{wrist}$ and $cam_{static}$, we constructed point clouds of the scene, which were used for occupancy (OCC) supervised training to enable the model to learn 3D geometric structures. The system operates on a ROS1-based communication framework to enable efficient interaction between the Nvidia 3090 server and the robotic arm.
In Figure 8, we design ten real-world tasks grouped into three difficulty levels. Easy: pick or push apple, pick banana or Coke bottle. Medium: open drawer, place lid on pot, pour Coke into cup. Hard: open drawer and place object inside, group similar items, store plush toys. We collect 100 teleoperation demonstrations per task (1000 total) for training. For evaluation, each task undergoes 10 trials with manual resets and clear success criteria (e.g., lifting an apple at least 5 cm). More details are provided in the supplementary due to space limits. We compare two baseline models: RoboFlamingo (2D) and RoboUniView (3D), which, like RoboTron-Mani, are both 3D models.
Figure 8: Real-world setup with 10 tasks, featuring a fixed Dalu base, UR3 arm, Robotiq gripper, wrist-mounted Intel D435, and ground-mounted Orbbec Gemini Pro. Hand-eye calibration and point cloud fusion enable OCC-supervised training. The system runs on ROS1 with an NVIDIA RTX 3090 GPU.
In Table 4 Right, RoboTron-Mani outperforms baselines across all difficulty levels, especially on medium and hard tasks, demonstrating stronger generalization and 3D reasoning. While RoboUniView performs reasonably on easy tasks, it struggles in more complex scenarios. RoboFlamingo, lacking explicit 3D scene modeling, shows weaker overall performance.
Model Easy Medium Hard
RoboTron-Mani 82.5% 70.0% 46.6%
RoboUniView 75.0% 36.6% 23.3%
RoboFlamingo 35.0% 6.6% 6.6%
Table 4: Success rates by task difficulty, indicates reproduced results.
8 RoboTron-Mani detailed information
@@ 10.2 Meta-World Dataset
Figure 9: Evaluation Datasets. We evaluate RoboTron-Mani across five simulation benchmarks and present policy rollout visualizations of the experiments. From left to right, the benchmarks include CALVIN, Meta-World, LIBERO, RoboCasa, and Robomimic. Experiment details can be found in Section 4.1.
@@ 10.9 Colosseum Dataset
11 Comparison with OpenVLA
To evaluate the superiority of our model architecture, we compare RoboTron-Mani with the currently best-performing OpenVLA. To ensure a fair comparison, we set the window size to 1, and train RoboTron-Mani from scratch, while OpenVLA is fine-tuned on the officially released weights.
As shown in Table 6, RoboTron-Mani outperforms OpenVLA (LoRA) in multiple metrics, particularly in Task 3-5 and average sequence length. This indicates that RoboTron-Mani can capture longer dependencies when handling tasks, thereby improving model accuracy. These results not only demonstrate the superior performance of the RoboTron-Mani architecture but also provide valuable references for future research.
Task Completed in a Sequence Avg Len 1 2 3 4 5 OpenVLA (LoRA) 78% 55% 29% 17% 8% 1.86 RoboTron-Mani(ours) 81% 54% 37% 25% 16% 2.15
Table 6: Performance comparison with OpenVLA on CALVIN.
@@ 12 Experiment Details
The success rates of the expert models in the Table 1 are organized from the following sources: The evaluation methods on the CALVIN [43] dataset are sourced from the official CALVIN leaderboard (url: http://calvin.cs.uni-freiburg.de/). In the Meta-World [68] dataset, the results of PAD [24], GR-1 [67], SuSIE [7], RT-2* [9], and RT-1 [8] come from PAD [24], while the results of PRISE [72] are derived from related papers. In the Libero [36] dataset, the results of QueST [47], VQ-BeT [32], PRISE [72], DiffusionPolicy [16], ACT [70], and ResNet-T [36] all come from QueST [47], while the results of MDT [58] and Distill-D [25] are sourced from MDT [58]; the results of MaIL [28], ATM [66], and MUTEX [59] come from their respective research papers. The results of RoboCasa [50] in the RoboCasa [50] dataset are sourced from related papers. In the Rt-1 [8] dataset, all results come from CogACT [34] paper.
The success rates of each task for RoboData on various datasets are shown in the Table 8,  9,  10,  11,  7.
Task Name Success Rate
CoffeePressButton 100%
CoffeeServeMug 50%
@@
CloseDrawer 100%
OpenDrawer 80%
TurnOffMicrowave 75%
TurnOnMicrowave 85%
PnPCabToCounter 45%
PnPCounterToCab 50%
PnPCounterToMicrowave 35%
PnPCounterToSink 40%
PnPCounterToStove 60%
PnPMicrowaveToCounter 70%
PnPSinkToCounter 40%
PnPStoveToCounter 40%
@@
Table 7: RoboTron-Mani Success Rates on Various Tasks in RoboCasa [50].
Task Success Rate
rotate_blue_block_right 82.6%
move_slider_right 98.2%
lift_red_block_slider 87.9%
place_in_slider 26.9%
turn_off_lightbulb 89.7%
turn_off_led 96.9%
push_into_drawer 68.1%
lift_blue_block_drawer 100.0%
lift_pink_block_slider 86.2%
open_drawer 93.1%
lift_pink_block_table 87.0%
turn_on_lightbulb 94.4%
rotate_blue_block_left 96.6%
push_blue_block_left 87.1%
close_drawer 91.0%
rotate_red_block_right 80.3%
push_pink_block_right 64.9%
push_red_block_right 59.7%
push_red_block_left 87.1%
lift_blue_block_table 87.2%
place_in_drawer 97.3%
move_slider_left 91.9%
rotate_red_block_left 84.6%
turn_on_led 93.0%
lift_red_block_table 93.0%
stack_block 55.4%
push_pink_block_left 91.2%
lift_blue_block_slider 85.2%
unstack_block 100.0%
rotate_pink_block_left 90.2%
lift_pink_block_drawer 85.7%
rotate_pink_block_right 63.5%
lift_red_block_drawer 93.3%
push_blue_block_right 48.4%
Table 8: RoboTron-Mani Success Rates on Various Tasks in CALVIN [43].
Task Success Rate
assembly-v2 100%
basketball-v2 100%
bin-picking-v2 70%
box-close-v2 85%
button-press-topdown-v2 100%
button-press-topdown-wall-v2 100%
button-press-v2 80%
button-press-wall-v2 85%
coffee-button-v2 90%
coffee-pull-v2 40%
coffee-push-v2 65%
dial-turn-v2 100%
disassemble-v2 80%
door-close-v2 100%
door-lock-v2 100%
door-open-v2 100%
door-unlock-v2 100%
hand-insert-v2 55%
drawer-close-v2 100%
drawer-open-v2 100%
faucet-open-v2 0%
faucet-close-v2 90%
hammer-v2 15%
handle-press-side-v2 100%
handle-press-v2 100%
handle-pull-side-v2 25%
handle-pull-v2 100%
lever-pull-v2 80%
peg-insert-side-v2 55%
pick-place-wall-v2 95%
pick-out-of-hole-v2 15%
reach-v2 75%
push-back-v2 100%
push-v2 90%
pick-place-v2 100%
plate-slide-v2 100%
plate-slide-side-v2 100%
plate-slide-back-v2 100%
plate-slide-back-side-v2 100%
peg-unplug-side-v2 25%
soccer-v2 20%
stick-push-v2 100%
stick-pull-v2 85%
push-wall-v2 100%
reach-wall-v2 85%
shelf-place-v2 45%
sweep-into-v2 95%
sweep-v2 100%
window-open-v2 60%
window-close-v2 100%
Table 9: RoboTron-Mani Success Rates on Various Tasks in Meta-World [68]
Method Pick Coke Can Move Near Open / Close Drawer Overall
RT-1-X 0.567 0.317 0.597 0.534
RT-2-X(55B) 0.787 0.779 0.250 0.607
Octo-Base 0.170 0.042 0.227 0.169
OpenVLA 0.163 0.462 0.356 0.248
HPT[1] 0.60 0.24 0.23 0.35
RoboTron-Mani 0.63 0.64 0.525 0.60
Table 10: SIMPLER evaluation results of different methods on RT-1. The “Overall” column reports the success rate averaged across the sub-tasks of all task types.
Task Index Success Rate Task Index Success Rate
0 100% 45 80%
1 85% 46 90%
2 95% 47 100%
3 95% 48 95%
4 85% 49 95%
5 60% 50 95%
6 100% 51 35%
7 100% 52 95%
8 90% 53 100%
9 80% 54 95%
10 100% 55 100%
11 95% 56 100%
12 90% 57 100%
13 75% 58 100%
14 95% 59 90%
15 95% 60 100%
16 100% 61 95%
17 95% 62 95%
18 95% 63 100%
19 100% 64 80%
20 100% 65 100%
21 85% 66 100%
22 90% 67 100%
23 70% 68 95%
24 100% 69 95%
25 100% 70 100%
26 90% 71 100%
27 75% 72 95%
28 100% 73 70%
29 100% 74 90%
30 100% 75 70%
31 100% 76 100%
32 50% 77 100%
33 85% 78 85%
34 100% 79 100%
35 85% 80 90%
36 90% 81 60%
37 100% 82 100%
38 80% 83 95%
39 85% 84 90%
40 95% 85 95%
41 100% 86 90%
42 100% 87 100%
43 100% 88 100%
44 95% 89 95%
Table 11: RoboTron-Mani Success Rates on Various Tasks in Libero [36].